OpenStudy (anonymous):

Let logb A=3; logb C=2; logb D=5 What is the value of logb A^5C^2/D^6?

OpenStudy (anonymous):

lobg A^5 = 5logb A = 15 logb C^2 = 2 logb C = 4 logb D^6 = 6 logb D = 30 can u continue from here - this is an exercise in applying laws of logs

OpenStudy (anonymous):

logb A^5 C^2 / D^6 = log A^5 + logb C^2 - logb D^6 so plug in the above values and you've got it